# MVPatch: More Vivid Patch for Adversarial Camouflaged Attacks on Object Detectors in the Physical World

### v2

@@ Abstract
Recent investigations demonstrate that adversarial patches can be utilized to manipulate the result of object detection models. However, the conspicuous patterns on these patches may draw more attention and raise suspicions among humans. Moreover, existing works have primarily focused on enhancing the efficacy of attacks in the physical domain, rather than seeking to optimize their stealth attributes and transferability potential. To address these issues, we introduce a dual-perception-based attack framework that generates an adversarial patch known as the More Vivid Patch (MVPatch). The framework consists of a model-perception degradation method and a human-perception improvement method. To derive the MVPatch, we formulate an iterative process that simultaneously constrains the efficacy of multiple object detectors and refines the visual correlation between the generated adversarial patch and a realistic image. Our method employs a model-perception-based approach that reduces the object confidence scores of several object detectors to boost the transferability of adversarial patches. Further, within the human-perception-based framework, we put forward a lightweight technique for visual similarity measurement that facilitates the development of inconspicuous and natural adversarial patches and eliminates the reliance on additional generative models. Additionally, we introduce the naturalness score and transferability score as metrics for an unbiased assessment of various adversarial patches’ natural appearance and transferability capacity. Extensive experiments demonstrate that the proposed MVPatch algorithm achieves superior attack transferability compared to similar algorithms in both digital and physical domains while also exhibiting a more natural appearance. These findings emphasize the remarkable stealthiness and transferability of the proposed MVPatch attack algorithm.
@@ I Introduction
Deep Neural Networks (DNNs) have achieved remarkable performance in manifold fields, such as computer vision[4, 5, 6, 7], natural language processing[8, 9], and automatic speech recognition[10, 11]. However, due to their lack of interpretability, DNNs are vulnerable to Adversarial Examples (AEs)[12, 13, 14], which raises concerns about their reliability in security-critical applications such as face recognition[15, 16] and autonomous driving[17, 18]. In general, AE attacks can be categorized into two types: Digital Attack (DA)[12, 13, 19, 20, 21, 22, 23, 24, 25, 26], which involves introducing digital perturbations into the input image to carry out the attack, and Physical Attack (PA)[27, 28, 29, 30, 31, 2, 32, 1, 33, 34, 3, 35, 16, 36, 37, 38], which directly targets real-world objects using adversarial perturbations, as illustrated in Fig.1(b) and Fig.1(a), respectively.
@@
Object Detection Models (ODMs)[39, 40, 41], comprised of deep neural layers, are extensively utilized in real-world applications such as people tracking[42], pedestrian and vehicle re-identification (Re-ID)[43, 44], and remote sensing[45]. Although ODMs enhance human convenience and improve the quality of life, they simultaneously pose a considerable risk to individual privacy and sensitive personal information[1, 3, 2, 32]. Furthermore, the capability of physical attacks to be transferred amplifies the vulnerability of ODMs[46, 47, 48]. In real-world scenarios, AEs could be often observed which are quasi-imperceptible to humans. For example, as depicted in Fig.1, an intruder appears under a surveillance camera holding an adversarial patch, which is camouflaged as Van Gogh’s painting, and the DNN-based object detection system will fail. It can even be said that the intruder is invisible under the camera, which will seriously threaten the public safety system Hence, it is imperative to conduct a thorough investigation into PA on ODMs.
Adversarial Patches (APs)[30] have emerged as a potent means to manipulate ODMs within the physical realm, offering several advantages such as input-independence and scene-independence, with substantial real-world impacts[22, 15, 49, 50, 3, 33]. However, much of the extant literature emphasizes enhancing the efficacy of the attack within this physical sphere, rather than seeking to optimize its stealth attributes and transferability potential. Hu et al.[1] considered how to improve the stealthiness of the adversarial patch by Generative Adversarial Network (GAN)[51]. In contrast, Guesmi et al.[38] proposed a lightweight approach to generate natural patches without GAN. Moreover, many scholars have endeavored to augment APs’ transferability. A case in point is the research by Wang et al.[46], who honed AP transferability within automated checkout systems through a bias-guided arrangement. In a similar endeavor, Wei et al.[47] devised a Pay No Attention (PNA) attack, wherein attention gradients are disregarded during backpropagation, thereby effectively improving backpropagation’s linearity. Furthermore, Huang et al.[48] modified a single surrogate model into a diverse ensemble structure, yielding more transferable adversarial examples. Nevertheless, numerous challenges persist, which include the substantial computational resource requirements associated with the use of GAN during the training process, and an unrelenting research focus on enhancing the stealthiness[2, 31, 52, 1, 32, 3, 35, 29, 53, 28, 38] and transferability[46, 47, 48] of APs. Significantly, there remains limited research simultaneously addressing both stealthiness and transferability issues related to APs, with an additional focus on computational resource efficiency.
To bridge this gap, we propose a dual-perception-based framework to generate a potent adversarial patch, referred to as the MVPatch (More Vivid Patch). This attack framework offers remarkable stealthiness and transferability capacities for attacking against ODMs. To improve the transferability potential, we present a model-perception degradation method for ensemble multiple object detectors achieved through the ensemble attack loss function. Additionally, to enhance the stealthiness attributes, we present a human-perception improvement method realized by the Compared Specified Image Similarity (CSS) loss function, which allows for the generation of natural and stealthy adversarial patches without the reliance on additional generative models. By integrating both the model-perception-based method and the human-perception-based method, we can generate a more powerful adversarial patch with significant transferable attack ability and invisibility. Moreover, we introduce the naturalness score and the transferability score as experimental evaluation metrics to assess the attack performance regarding stealthiness and transferability. We exploit the YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny as threat models and Faster RCNN, SSD, and YOLOv5 as transfer attack models. Based on extensive experiments conducted in both digital and physical scenarios, along with several independent subjective surveys, our proposed MVPatch consistently generates more natural and transferable results compared to several state-of-the-art baselines. We present the primary specialization of MVPatch, along with the distinguishing features of adversarial patches that possess meaning and those that are devoid of meaning. The main contributions of this work can be summarized as follows:
  • To the best of our knowledge, we are the first to propose a dual-perception-based attack framework for generating an adversarial patch (MVPatch). This patch offers both stealthiness and transferability by exploiting both model-perception-based and human-perception-based methods.
  • We present the model-perception-based method, an attack method that reduces the object confidence scores of an ensemble of multiple object detectors using an ensemble attack loss function. This enhances the transferability of the adversarial patches, significantly improving attack performance in both digital and physical domains compared to state-of-the-art baselines.
@@
  • We propose a human-perception-based method, a lightweight visual similarity measurement method that generates natural and stealthy adversarial patches without the need for additional generative models. This allows us to generate patches based on specific images.
@@
  • We employ naturalness and transferability scores as evaluation metrics to assess the naturalness and attack transferability of the diverse adversarial patches in our experiment.
@@
The remainder of this paper is organized as follows. Section II provides a concise literature review on adversarial examples and adversarial patches in various scenarios, encompassing both digital and physical contexts. In Section III, we introduce the definition and characteristics of adversarial patch attacks, along with detailed information regarding the input and output of the threat models. The generation process of our proposed MVPatch is then presented in this section. Section IV introduces the environment and dataset in the experiment, followed by the provision of relevant experimental details. Additionally, we propose evaluation metrics for the experiment. Section V presents empirical evidence that showcases the effectiveness of the proposed MVPatch through comprehensive evaluations conducted in both digital and physical environments. Ultimately, we summarize the paper and present the further research directions in Section VI.
@@ II Related Work
In this section, we briefly review the literature related to adversarial examples and adversarial patches in diverse scenarios, such as in the digital domains and physical domains. Furthermore, we discuss the recent issues to be solved and present the necessity of our work.
@@ II-A Digital Attack
Adversarial Examples (AEs) are elaborately designed perturbations that are imperceptible to humans but could mislead Deep Neural Networks (DNNs). In 2014, Szegedy et al.[12] discovered a special property of DNNs known as Adversarial Examples (AEs). They successfully attacked a DNN using the L-BFGS algorithm, causing it to produce incorrect predictions. In 2015, Goodfellow et al.[13] introduced the Fast Gradient Sign Method (FGSM), which improved the effectiveness and success rates of attacks on neural network classification models. Kurakin et al.[22] developed the Basic Iterative Method (BIM) and demonstrated the application of AEs in the physical world. Papernot et al.[23] focused on partial attack and presented the Jocobian-Based Saliency Map Attack (JSMA) algorithm. Madry et al.[19] proposed Project Gradient Descent (PGD), an innovative framework that combines attack and defense algorithms to enhance the robustness of DNNs. There are numerous varieties of adversarial attacks in the digital domain, e.g., Deepfool[21], C&W[20], ZOO[24], Universal Perturbation[25], One-Pixel attack[26] and some attacks on large language vision models[54, 55, 56].
@@ II-B Physical Attack
Recently, it has been observed that printed AEs can effectively deceive neural network models in the physical domain. Athalya et al.[27] developed robust three-dimensional AEs that can mitigate the impact of diverse angle transformations. Sharif et al.[28] successfully attacked facial recognition systems by constructing adversarial eyeglasses, while Komkov et al.[29] proposed the construction of an adversarial hat for the same purpose. In 2018, Brown et al.[30] introduced the concept of Adversarial Patch (AP). Liu et al.[31] proposed DPatch to deceive object detection models, while at around the same time, Thy et al.[2] introduced AdvPatch to mislead automated surveillance systems. Wu et al. [32] presented the AdvCloack to make humans invisible in the object detector. Hu et al.[1] designed a more natural patch based on the AdvPatch[2]. Xu et al.[33] proposed to construct an adversarial T-shirt to mislead the object detection systems. Huang et al.[34] introduced the Transfer-based Self-ensemble Attack (T-SEA) for attacking object detectors using ensemble attacks, effectively improving the transferable attack performance. Hu et al.[3] proposed the Toroidal-Cropping-based Expandable Generative Attack (TC-EGA) algorithm, which applies adversarial textures to clothing to improve the stealthiness of the adversarial patch. Wang et al.[35] constructed the adversarial patch based on the attention mechanism for the car camouflage. Wei et al.[16] investigated methods to enhance the attack performance of adversarial patches by focusing on their spatial positioning. Liu et al.[36] proposed a Perceptual-Sensitive Generative Adversarial Network (PS-GAN) to enhance the visual fidelity and attacking performance for the adversarial patches. Xue et al.[37] presented the Diffusion-Based Projected Gradient Desent (DIff-PGD) to generate realistic adversarial examples by using diffusion models. Guesmi et al.[38] proposed AdvART to attack object detectors by a simple method without generative models.
Despite the significant results achieved by prior research on adversarial patches in the physical world, there are still some issues to be addressed, such as the trade-off between transferability[34, 1, 57] and stealthiniess[27, 29, 32, 33, 3, 35, 58, 36, 37, 38] of the adversarial patch. Furthermore, most recent research focuses on generating meaningless adversarial patches. The meaningless adversarial patch means that has no sense for human eye perception as Fig.1(d) illustrates. To the best of our knowledge, only one study has systematically examined the investigation of meaningful and transferable adversarial patches through the utilization of a generative adversarial network (GAN) by Hu et al.[1]. However, training a generative adversarial network is challenging and computationally expensive. To address these issues, we propose the MVPatch algorithm, which outperforms other state-of-the-art baseline methods in terms of transferable and stealthy attacks both in the digital and physical scenarios while maintaining a lightweight approach.

### v3

@@ Abstract
Recent studies have shown that Adversarial Patches (APs) can effectively manipulate object detection models. However, the conspicuous patterns often associated with these patches tend to attract human attention, posing a significant challenge. Existing research has primarily focused on enhancing attack efficacy in the physical domain while often neglecting the optimization of stealthiness and transferability. Furthermore, applying APs in real-world scenarios faces major challenges related to transferability, stealthiness, and practicality. To address these challenges, we introduce generalization theory into the context of APs, enabling our iterative process to simultaneously enhance transferability and refine visual correlation with realistic images. We propose a Dual-Perception-Based Framework (DPBF) to generate the More Vivid Patch (MVPatch), which enhances transferability, stealthiness, and practicality. The DPBF integrates two key components: the Model-Perception-Based Module (MPBM) and the Human-Perception-Based Module (HPBM), along with regularization terms. The MPBM employs ensemble strategy to reduce object confidence scores across multiple detectors, thereby improving AP transferability with robust theoretical support. Concurrently, the HPBM introduces a lightweight method for achieving visual similarity, creating natural and inconspicuous adversarial patches without relying on additional generative models. The regularization terms further enhance the practicality of the generated APs in the physical domain. Additionally, we introduce naturalness and transferability scores to provide an unbiased assessment of APs. Extensive experimental validation demonstrates that MVPatch achieves superior transferability and a natural appearance in both digital and physical domains, underscoring its effectiveness and stealthiness.
@@ I Introduction
Deep Neural Networks (DNNs) have achieved remarkable performance in manifold fields, such as computer vision [4, 5, 6, 7], natural language processing[8, 9], and automatic speech recognition [10, 11]. However, due to their lack of interpretability, DNNs are vulnerable to Adversarial Examples (AEs) [12, 13, 14, 15], which raises concerns about their reliability in security-critical applications such as face recognition [16, 17] and autonomous driving [18, 19]. In general, AE attacks can be categorized into two types: Digital Attack (DA) [12, 13, 20, 21, 22, 23, 24, 25, 26, 27], which involves introducing digital perturbations into the input image to carry out the attack, and Physical Attack (PA) [28, 29, 30, 31, 32, 2, 33, 1, 34, 35, 3, 36, 17, 37, 38, 39], which directly targets real-world objects using adversarial perturbations, as illustrated in Fig. 1(b) and Fig. 1(a), respectively.
@@
Object Detection Models (ODMs) [40, 41, 42], comprised of deep neural layers, are extensively utilized in real-world applications such as people tracking [43], pedestrian re-identification (Re-ID) [44], and remote sensing [45]. Although ODMs enhance human convenience and improve the quality of life, they simultaneously pose a considerable risk to individual privacy and sensitive personal information [1, 3, 2, 33]. Furthermore, the capability of physical attacks to be transferred amplifies the vulnerability of ODMs [46, 47, 48, 49, 50, 51]. In real-world scenarios, AEs could be often observed which are quasi-imperceptible to humans. For example, as depicted in Fig. 1, an intruder appears under a surveillance camera holding an adversarial patch, which is camouflaged as Van Gogh’s painting, and the DNN-based object detection system will fail. It can even be said that the intruder is invisible under the camera, which will seriously threaten the public safety system Hence, it is imperative to conduct a thorough investigation into PA on ODMs.
Adversarial Patches (APs) [31] have emerged as a potent means to manipulate ODMs within the physical realm, offering several advantages such as input-independence and scene-independence, with substantial real-world impacts [23, 16, 52, 53, 3, 34]. However, despite their potential, adversarial patches face three major issues that hinder their broader applicability and effectiveness: Transferability, Stealthiness, and Practicality. Transferability can be enhanced through methods like gradient-based optimization (e.g., MI [54], NI [55]), input transformation (e.g., DI [56], SI [55]), and model ensemble methods (e.g., MI [54], SVRE [57]), with the latter being the most effective yet not fully understood. Stealthiness involves generating realistic APs using GAN-based methods (e.g., Natural Patch [1]), Diffusion-based methods (e.g., Diff-PGD [38]), and Regularization-based methods (e.g., AdvART [39], DAP [58]), each with trade-offs in computational cost and efficiency. Practicality concerns the transition from digital to physical domains [30, 29, 28, 35], where APs face challenges like performance degradation due to varying distances and angles. Moreover, much of the existing literature either focuses on enhancing the transferability of adversarial patches [46, 47, 48, 59] while overlooking their stealthiness, or prioritizes stealthiness [2, 32, 60, 1, 33, 3, 36, 30, 61, 29, 39] at the expense of transferability. Notably, there is limited research that simultaneously addresses both stealthiness and transferability, while also considering computational resource efficiency.
To address these challenges, we propose a Dual-Perception-Based Framework (DPBF) for generating a powerful adversarial patch, termed the More Vivid Patch (MVPatch). This framework enhances the transferability, stealthiness, and practicality of adversarial patches for attacking object detection models (ODMs). To improve transferability, we introduce the Model-Perception-Based Module (MPBM), which employs an ensemble strategy with multiple object detectors. The MPBM leverages the ensemble strategy to boost adversarial attack transferability, supported by theoretical insights from generalization theory. For enhanced stealthiness, we present the Human-Perception-Based Module (HPBM). This module generates adversarial patches that closely mimic a specified image in Hilbert space, using diverse transformations and the Frobenius norm to minimize generalization error. This approach ensures the patches are natural and inconspicuous without relying on additional generative models. To improve real-world practicality, we incorporate regularization terms such as Total Variation (TV) and Non-Printable Score (NPS) into the DPBF. By integrating the MPBM, HPBM, and regularization terms, we create a more effective adversarial patch with significant transferability and stealthiness while maintaining practicality. Additionally, we introduce naturalness and transferability scores as experimental metrics to assess the performance of the adversarial patches. We evaluate these patches using threat models including YOLOv2, YOLOv3, YOLOv3-tiny, YOLOv4, and YOLOv4-tiny, as well as transfer attack models such as Faster R-CNN, SSD, and YOLOv5. Extensive experiments in both digital and physical scenarios, along with independent subjective surveys, demonstrate that our MVPatch consistently outperforms several state-of-the-art baselines in terms of naturalness and transferability.
The main contributions can be summarized as follows:
  • To the best of our knowledge, we are the first to introduce generalization theory to the field of adversarial patches. This approach provides robust theoretical support for generating patches that maintain transferability, stealthiness, and practicality through our Dual-Perception-Based Framework (DPBF).
  • We present the Model-Perception-Based Module (MPBM), which reduces object confidence scores through an ensemble strategy. Utilizing generalization theory, the MPBM demonstrates that this ensemble approach enhances both generalization and stability, thereby improving the transferability of adversarial patches (APs) compared to single-model strategies.
@@
  • We propose a Human-Perception-Based Module (HPBM) that generates natural and stealthy adversarial patches without the need for additional generative models, while ensuring minimal generalization error.
@@
  • We employ naturalness and transferability scores as evaluation metrics to assess the naturalness and attack transferability of diverse adversarial patches in our experiments.
@@
The remainder of this paper is organized as follows: Section II reviews the literature on adversarial examples, covering both digital and physical domains, and explores generalization aspects for adversarial attacks. Section III defines adversarial patch attacks, discusses their characteristics, and details the input and output of the threat models, as well as the generation process of the proposed MVPatch. Section IV describes the experimental environment and dataset, including the experimental setup and evaluation metrics. Section V demonstrates the empirical effectiveness of MVPatch through extensive evaluations in both digital and physical settings. Finally, Section VI summarizes the findings, discusses limitations, and outlines future research directions.
@@ II-A Digital Attack
Adversarial Examples (AEs) are perturbations that are imperceptible to humans but can mislead Deep Neural Networks (DNNs). In 2014, Szegedy et al. [12] first discovered AEs, successfully attacking a DNN using the L-BFGS algorithm. In 2015, Goodfellow et al. [13] introduced the Fast Gradient Sign Method (FGSM), enhancing attack effectiveness on neural network classifiers. Kurakin et al. [23] developed the Basic Iterative Method (BIM) and demonstrated AEs’ applicability in the physical world. Papernot et al. [24] presented the Jocobian-Based Saliency Map Attack (JSMA) for partial attacks. Madry et al. [20] proposed Project Gradient Descent (PGD) to enhance DNN robustness through a combined attack and defense framework. Other notable digital adversarial attacks include Deepfool [22], C&W [21], ZOO [25], Universal Perturbation [26], One-Pixel attack [27], and attacks on hasing retrival [62] and large language vision models [63, 64, 65].
@@ II-B Physical Attack
Recently, it has been observed that printed Adversarial Examples (AEs) can effectively deceive neural network models in the physical domain. Athalya et al. [28] developed robust 3D AEs mitigating diverse angle transformations. Sharif et al. [29] and Komkov et al. [30] attacked facial recognition systems with adversarial eyeglasses and hats, respectively. Brown et al. [31] introduced the concept of Adversarial Patches (APs). Liu et al. [32] proposed DPatch for object detection models, and Thy et al. [2] introduced AdvPatch for misleading surveillance systems. Wu et al. [33] presented AdvCloak to make humans invisible to object detectors. Hu et al. [1] designed a more natural patch based on AdvPatch, and Xu et al. [34] created an adversarial T-shirt for object detection. Huang et al. [35] introduced T-SEA for ensemble attacks on object detectors. Hu et al. [3] proposed TC-EGA, applying adversarial textures to clothing for stealthiness. Wang et al. [36] developed an adversarial patch for car camouflage using attention mechanisms. Wei et al. [17] focused on enhancing attack performance via spatial positioning of patches. Liu et al. [37] proposed PS-GAN for improved visual fidelity and attack performance. Xue et al. [38] introduced DIff-PGD to generate realistic AEs using diffusion models. Guesmi et al. [39] proposed AdvART for attacking object detectors without generative models.
II-C Generalization Analysis for Adversarial Patch Attack
Generalization evaluates how well a model trained on a dataset performs on unseen data, typically using Empirical Risk Minimization (ERM) [66]. Given data $\{(x_{i},y_{i}),i=1,\dots,n\}$ from a probability distribution $\mathcal{D}$ on $\Omega\times\{-1,1\}$, a class of functions $\mathcal{F}:\Omega\rightarrow\mathbb{R}$ and and a loss function $l$, ERM finds a minimizer of the empirical loss:
$f^{*}=\operatorname*{arg\,min}_{f\in\mathcal{F}}\mathcal{L}_{\text{emp}}(f):=\operatorname*{arg\,min}_{f\in\mathcal{F}}\sum_{i}l(f(x_{i}),y_{i}).$ (1)
The complexity of the model $\mathcal{F}$ can be assessed using approaches such as VC-dimension [67], Rademacher complexity [68], and covering numbers [69]. These analyses typically provide bounds on the generalization gap, which is the difference between the empirical loss and the expected loss:
$|\left[l(f^{*}(x),y)\right]-\mathcal{L}_{\text{emp}}(f^{*}(x))|<O^{*}(\sqrt{\frac{c}{n}}).$ (2)
From Eq. 2, we observe that the generalization error depends on both model complexity $c$ and the number of samples $n$. Recent studies have investigated the relationship between adversarial robustness and generalization [70, 71]. Huang et al. [35] introduced T-SEA, which utilizes the insights from Eq. 2 to improve the transferability of generated adversarial patches.
Despite significant advancements in adversarial patches for the physical world, challenges persist, particularly in balancing transferability [35, 1, 72] and stealthiness [28, 30, 33, 34, 3, 36, 73, 37, 38, 39]. Recent research often focuses on generating meaningless adversarial patches that lack visual coherence for human perception, as shown in Fig. 1(d). Hu et al. [1] systematically explored meaningful and transferable adversarial patches using GANs, although training GANs is computationally expensive. Moreover, there is a lack of robust theoretical support for addressing the issues of transferability, stealthiness, and practicality. To address these challenges, we propose a Dual-Perception-Based Framework (DPBF) to generate the More Vivid Patch (MVPatch) with enhanced transferability, stealthiness, and practicality. Inspired by T-SEA [35], we incorporate generalization theory to comprehensively analyze and improve the generalization and stability of adversarial patches, enhancing both their transferability and stealthiness. Our MVPatch outperforms state-of-the-art methods in both digital and physical scenarios, achieving superior transferability and stealthiness with a lightweight approach.